OpenStudy (anonymous):

Find the other endpoint of the line segment with the given endpoint and mid point. 1. Endpoint:(-1, 9), midpoint: (-9, -10) 2. Endpoint: (2, 5), midpoint: (5, 1) 3. Endpoint:(9, -10), midpoint: (4, 8) Critical Thinking 4. Find the point that is one-fourth of the way from (2, 4) to (10, 8).

OpenStudy (anonymous):

For x: (-1+x)/2 = -9 -1/2 + x/2 = -18/2 x/2 = -17/2 x=-17 For y: (9+y)/2 = -10 9/2 +y/2 = -20/2 y/2 = -29/2 y = -29 Your point: (-17,-29)

OpenStudy (anonymous):

OH! Here we go! It's simpler than I though. Sorry about that. http://www.algebra.com/algebra/homework/Coordinate-system/Coordinate-system.faq.question.176789.html

OpenStudy (anonymous):

*thought

OpenStudy (anonymous):

Diff in x: 10-2 = 8 Diff in y: 8-4=4 1/4 of x: (1/4)* 8 = 2 1/4 of y: (1/4)*4 = 1 original points: (2,4) End Point = (x1 + diff in x, y1 + diff in y) End point = (2 + 2, 4 + 1) = (4,5)

OpenStudy (anonymous):

The distance is correct if you check (2,4) and (10,8) vs. (2,4) and (4,5) b/c the second is 1/4 of the distance of the first one.

OpenStudy (anonymous):

@purdrew is number 4 making sense? B/c this looks like the "hardest" of them all.